An off-ramp is graded at a slope of 4.00 degrees with the horizontal. How many meters of pavement will be needed if the ramp takes vehicles down a vertical descent of 120.0 m?
1
Expert's answer
2012-06-01T09:08:06-0400
x = 120.0 / cos(90.00 - 4.00) = 120.0 / cos(86.00) = 1720.27 m